If cotA=7/8 then evaluate
1) (1+sinA)(1-sinA)/(1+cosA)(1-cosA)
2) cot^2A

Answer by lwsshak3(11628)

If cotA=7/8 then evaluate
1) (1+sinA)(1-sinA)/(1+cosA)(1-cosA)
2) cot^2A
***
use x for A
cotx=7/8
1) (1+sinx)(1-sinx)/(1+cosx)(1-cosx)
1-sin^2(x)/1-cos^2(x)
cos^2(x)/sin^2(x)=cot^2(x)=49/64
..
2) cot^2x=49/64